OpenStudy (anonymous):

Find X intercepts x^2 (25-x^2)^1/2

OpenStudy (anonymous):

what are the A B C D

OpenStudy (anonymous):

no plug in 0 for y the awnser is 5,0 and -5,0

OpenStudy (anonymous):

\[x^{2}\]

OpenStudy (anonymous):

i need to know how to get it

OpenStudy (anonymous):

someone show me how to get it

OpenStudy (anonymous):

set equal to zero \[\Large x^2 \sqrt{25-x^2}=0\] now solve \[\Large x^2=0\] \[\Large \sqrt{25-x^2}=0\]

OpenStudy (anonymous):

how do u get 5,0 and -5,0

OpenStudy (anonymous):

for x^2=0 means x=0,0 now solve the following first square of both sides \[\sqrt{25-x^2}=0\]

OpenStudy (anonymous):

the awnsers are y int 0,0 x intercepts 5 and -5.

OpenStudy (anonymous):

how u get them

OpenStudy (anonymous):

why are you keep on insisting the answer ? Just follow the steps given above sorry OS do not Provide answers. if you want to understand topic and study it . you are Most welcome.

OpenStudy (anonymous):

no i got awnsers i just dont understand how to do it

OpenStudy (anonymous):

let me give you an Example if you have to solve \[\sqrt{36-x^2}=0\] take square of both sides \[\Large \implies 36-x^2=0\] or \[\Large x^2=36\] take square root of both sides \[\Large x=\pm 6\] does this help ?

OpenStudy (anonymous):

i got it but i hhave another one if u can help

OpenStudy (anonymous):

first Close this Question .post it in Question box will help you